$SPAD/src/input unittest1.input

Timothy Daly December 3, 2016

Abstract

Unit test the user level commands

Contents

2

— * —

)set break resume
)sys rm -f unittest1.output
)spool unittest1.output
)set mes auto off
)clear all
--S 1 of 97
)with API
--R   )library cannot find the file API.
--E 1
--S 2 of 97 this command generates random output, ok
--)apropos matrix
--E 2
--S 3 of 97
)what categories set
--R
--R------------------------------- Categories --------------------------------
--R
--RCategories with names matching patterns:
--R     set
--R
--R CACHSET  CachableSet                  FSAGG    FiniteSetAggregate
--R MSETAGG  MultisetAggregate
--R NTSCAT   NormalizedTriangularSetCategory
--R OMSAGG   OrderedMultisetAggregate     ORDSET   OrderedSet
--R PSETCAT  PolynomialSetCategory        RSETCAT  RegularTriangularSetCategory
--R SETAGG   SetAggregate                 SETCAT   SetCategory
--R SETCATD  SetCategoryWithDegree
--R SFRTCAT  SquareFreeRegularTriangularSetCategory
--R SNTSCAT  SquareFreeNormalizedTriangularSetCategory
--R TSETCAT  TriangularSetCategory
--E 3
--S 4 of 97
)what commands set
--R
--R--------------- System Commands for User Level: development ---------------
--R
--RSystem commands at this level matching patterns:
--R     set
--R
--Rset
--R
--E 4

--S 5 of 97 this command generates random output ok
--)what domains set
--E 5
--S 6 of 97 this command generates random output ok
--)what operations set
--E 6
--S 7 of 97
)what packages set
--R
--R-------------------------------- Packages ---------------------------------
--R
--RPackages with names matching patterns:
--R     set
--R
--R AFALGGRO AffineAlgebraicSetComputeWithGroebnerBasis
--R AFALGRES AffineAlgebraicSetComputeWithResultant
--R FSAGG2   FiniteSetAggregateFunctions2 LAZM3PK  LazardSetSolvingPackage
--R PRJALGPK ProjectiveAlgebraicSetPackage
--R PSETPK   PolynomialSetUtilitiesPackage
--R QALGSET2 QuasiAlgebraicSet2
--R RSDCMPK  RegularSetDecompositionPackage
--R RSETGCD  RegularTriangularSetGcdPackage
--R SFRGCD   SquareFreeRegularTriangularSetGcdPackage
--R SRDCMPK  SquareFreeRegularSetDecompositionPackage
--E 7
--S 8 of 97
)what synonym set
--R
--R------------------------- System Command Synonyms -------------------------
--R
--R   No user-defined synonyms satisfying patterns:
--R       set
--R
--E 8
--S 9 of 97 this command generates random output ok
--)what things set
--E 9
--S 10 of 97 this command generates random output ok
--)apropos set
--E 10
--S 11 of 97
)prompt
--R---------------------------- The prompt Option ----------------------------
--R

--R Description: set type of input prompt to display
--R
--R The prompt option may be followed by any one of the following:
--R
--R    none
--R    frame
--R    plain
--R -> step
--R    verbose
--R
--R The current setting is indicated.
--R
--E 11
--S 12 of 97
)version
--R
--IValue = "Saturday February 21, 2009 at 17:59:27 "
--E 12
--S 13 of 97
)zsys )from )c
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 13
--S 14 of 97
)zsys )from )d
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 14
--S 15 of 97
)zsys )from )dt
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...

--R
--E 15
--S 16 of 97
)zsys )from )ct
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 16
--S 17 of 97
)zsys )from )ctl
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 17
--S 18 of 97
)zsys )from )ec
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 18
--S 19 of 97
)zsys )from )ect
--R
--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 19
--S 20 of 97
)zsys )from )e
--R

--R
--R   >> System error:
--I   Cannot open the file /research/test/mnt/fedora10/../../src/interp/TAGS.
--R
--R   Continuing to read the file...
--R
--E 20
--S 21 of 97
)zsys )from )version
--R
--E 21
--S 22 of 97
)zsys )from )update
--R
--R
--R   >> System error:
--R   /UPDATE-1 [or a callee] requires more than one argument.
--R
--R   Continuing to read the file...
--R
--E 22
--S 23 of 97
)zsys )from )patch
--R
--R
--R   >> System error:
--R   The function /UPDATE-LIB-1 is undefined.
--R
--R   Continuing to read the file...
--R
--E 23
--S 24 of 97
)zsys )from )there 1
--R
--R
--R   Unknown option: there
--R   Available options are c ct e ec ect cls pause update patch compare record
--R
--E 24
--S 25 of 97
)zsys )from )compare
--R
--R   An argument is required for compare
--E 25

--S 26 of 97
)zsys )from )record
--R
--R   An argument is required for record
--E 26
--S 27 of 97
)summary
--R
--E 27
--S 28 of 97
--R)credits
--RAn alphabetical listing of contributors to AXIOM:
--RCyril Alberga          Roy Adler              Christian Aistleitner
--RRichard Anderson       George Andrews         S.J. Atkins
--RHenry Baker            Martin Baker           Stephen Balzac
--RYurij Baransky         David R. Barton        Gerald Baumgartner
--RGilbert Baumslag       Michael Becker         Nelson H. F. Beebe
--RJay Belanger           David Bindel           Fred Blair
--RVladimir Bondarenko    Mark Botch             Alexandre Bouyer
--RPeter A. Broadbery     Martin Brock           Manuel Bronstein
--RStephen Buchwald       Florian Bundschuh      Luanne Burns
--RWilliam Burge
--RQuentin Carpent        Robert Caviness        Bruce Char
--ROndrej Certik          Cheekai Chin           David V. Chudnovsky
--RGregory V. Chudnovsky  Josh Cohen             Christophe Conil
--RDon Coppersmith        George Corliss         Robert Corless
--RGary Cornell           Meino Cramer           Claire Di Crescenzo
--RDavid Cyganski
--RNathaniel Daly         Timothy Daly Sr.       Timothy Daly Jr.
--RJames H. Davenport     Didier Deshommes       Michael Dewar
--RJean Della Dora        Gabriel Dos Reis       Claire DiCrescendo
--RSam Dooley             Lionel Ducos           Lee Duhem
--RMartin Dunstan         Brian Dupee            Dominique Duval
--RRobert Edwards         Heow Eide-Goodman      Lars Erickson
--RRichard Fateman        Bertfried Fauser       Stuart Feldman
--RJohn Fletcher          Brian Ford             Albrecht Fortenbacher
--RGeorge Frances         Constantine Frangos    Timothy Freeman
--RKorrinn Fu
--RMarc Gaetano           Rudiger Gebauer        Kathy Gerber
--RPatricia Gianni        Samantha Goldrich      Holger Gollan
--RTeresa Gomez-Diaz      Laureano Gonzalez-Vega Stephen Gortler
--RJohannes Grabmeier     Matt Grayson           Klaus Ebbe Grue
--RJames Griesmer         Vladimir Grinberg      Oswald Gschnitzer
--RJocelyn Guidry
--RGaetan Hache           Steve Hague            Satoshi Hamaguchi
--RMike Hansen            Richard Harke          Bill Hart
--RVilya Harvey           Martin Hassner         Arthur S. Hathaway
--RDan Hatton             Waldek Hebisch         Karl Hegbloom

--RRalf Hemmecke          Henderson              Antoine Hersen
--RGernot Hueber
--RPietro Iglio
--RAlejandro Jakubi       Richard Jenks
--RKai Kaminski           Grant Keady            Tony Kennedy
--RTed Kosan              Paul Kosinski          Klaus Kusche
--RBernhard Kutzler
--RTim Lahey              Larry Lambe            Kaj Laurson
--RFranz Lehner           Frederic Lehobey       Michel Levaud
--RHoward Levy            Liu Xiaojun            Rudiger Loos
--RMichael Lucks          Richard Luczak
--RCamm Maguire           Francois Maltey        Alasdair McAndrew
--RBob McElrath           Michael McGettrick     Ian Meikle
--RDavid Mentre           Victor S. Miller       Gerard Milmeister
--RMohammed Mobarak       H. Michael Moeller     Michael Monagan
--RMarc Moreno-Maza       Scott Morrison         Joel Moses
--RMark Murray
--RWilliam Naylor         Patrice Naudin         C. Andrew Neff
--RJohn Nelder            Godfrey Nolan          Arthur Norman
--RJinzhong Niu
--RMichael O'Connor       Summat Oemrawsingh     Kostas Oikonomou
--RHumberto Ortiz-Zuazaga
--RJulian A. Padget       Bill Page              David Parnas
--RSusan Pelzel           Michel Petitot         Didier Pinchon
--RAyal Pinkus            Jose Alfredo Portes
--RClaude Quitte
--RArthur C. Ralfs        Norman Ramsey          Anatoly Raportirenko
--RAlbert D. Rich         Michael Richardson     Renaud Rioboo
--RJean Rivlin            Nicolas Robidoux       Simon Robinson
--RRaymond Rogers         Michael Rothstein      Martin Rubey
--RPhilip Santas          Alfred Scheerhorn      William Schelter
--RGerhard Schneider      Martin Schoenert       Marshall Schor
--RFrithjof Schulze       Fritz Schwarz          Steven Segletes
--RNick Simicich          William Sit            Elena Smirnova
--RJonathan Steinbach     Fabio Stumbo           Christine Sundaresan
--RRobert Sutor           Moss E. Sweedler       Eugene Surowitz
--RMax Tegmark            James Thatcher         Balbir Thomas
--RMike Thomas            Dylan Thurston         Steve Toleque
--RBarry Trager           Themos T. Tsikas
--RGregory Vanuxem
--RBernhard Wall          Stephen Watt           Jaap Weel
--RJuergen Weiss          M. Weller              Mark Wegman
--RJames Wen              Thorsten Werther       Michael Wester
--RJohn M. Wiley          Berhard Will           Clifton J. Williamson
--RStephen Wilson         Shmuel Winograd        Robert Wisbauer
--RSandra Wityak          Waldemar Wiwianka      Knut Wolf
--RClifford Yapp          David Yun
--RVadim Zhytnikov        Richard Zippel         Evelyn Zoernack
--RBruno Zuercher         Dan Zwillinger
--E 28

--S 29 of 97
)set expose
--R
--R---------------------------- The expose Option ----------------------------
--R
--R Description: control interpreter constructor exposure
--R
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   basic
--R                                categories
--R                                  naglink
--R                                   anna
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   The following constructors are explicitly hidden in the current
--R      frame:
--R                there are no explicitly hidden constructors
--R
--E 29
--S 30 of 97
)set expose add
--R----------------------------- The add Option ------------------------------
--R   The following groups are explicitly exposed in the current frame
--I      (called frame0 ):
--R                                   basic
--R                                categories
--R                                  naglink
--R                                   anna
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   When )set expose add is followed by no arguments, the information
--R      you now see is displayed. The arguments group and constructor are
--R      used to specify exposure groups or an explicit constructor to be
--R      added to the local frame exposure data. Issue
--R                            )set expose add group
--R      or
--R                         )set expose add constructor
--R      for more information.
--E 30
--S 31 of 97

)set expose drop
--R----------------------------- The drop Option -----------------------------
--R   The following constructors are explicitly hidden in the current
--R      frame:
--R                there are no explicitly hidden constructors
--R
--R   When )set expose drop is followed by no arguments, the information
--R      you now see is displayed. The arguments group and constructor are
--R      used to specify exposure groups or an explicit constructor to be
--R      dropped from the local frame exposure data. Issue
--R                           )set expose drop group
--R      or
--R                        )set expose drop constructor
--R      for more information.
--E 31
--S 32 of 97
)set expose add group
--R
--R---------------------------- The group Option -----------------------------
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   basic
--R                                categories
--R                                  naglink
--R                                   anna
--R
--Rbasic         naglink       anna          categories    Hidden
--Rdefaults
--E 32
--S 33 of 97
)set expose add constructor
--R------------------------- The constructor Option --------------------------
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--E 33
--S 34 of 97
)set expose drop group
--R---------------------------- The group Option -----------------------------
--R   When followed by one or more exposure group names, this option
--R      allows you to remove those groups from the local frame exposure
--R      data.
--R
--R   The following groups are explicitly exposed in the current frame
--I      (called frame0 ):
--R                                   basic
--R                                categories

--R                                  naglink
--R                                   anna
--E 34
--S 35 of 97
)set expose drop constructor
--R------------------------- The constructor Option --------------------------
--R   When followed by one or more constructor names, this option allows
--R      you to explicitly hide constructors in this frame.
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   The following constructors are explicitly hidden in the current
--R      frame:
--R                there are no explicitly hidden constructors
--E 35
--S 36 of 97
)show SQMATRIX
--R
--R SquareMatrix(ndim: NonNegativeInteger,R: Ring)  is a domain constructor
--R Abbreviation for SquareMatrix is SQMATRIX
--R This constructor is not exposed in this frame.
--R Issue )edit bookvol10.3.pamphlet to see algebra source code for SQMATRIX
--R
--R------------------------------- Operations --------------------------------
--R ?*? : (R,%) -> %                      ?*? : (%,R) -> %
--R ?*? : (%,%) -> %                      ?*? : (Integer,%) -> %
--R ?*? : (PositiveInteger,%) -> %        ?**? : (%,PositiveInteger) -> %
--R ?+? : (%,%) -> %                      ?-? : (%,%) -> %
--R -? : % -> %                           ?=? : (%,%) -> Boolean
--R D : % -> % if R has DIFRING           D : (%,(R -> R)) -> %
--R 1 : () -> %                           0 : () -> %
--R ?^? : (%,PositiveInteger) -> %        antisymmetric? : % -> Boolean
--R coerce : % -> Matrix(R)               coerce : R -> %
--R coerce : Integer -> %                 coerce : % -> OutputForm
--R copy : % -> %                         diagonal? : % -> Boolean
--R diagonalMatrix : List(R) -> %         diagonalProduct : % -> R
--R elt : (%,Integer,Integer) -> R        elt : (%,Integer,Integer,R) -> R
--R empty : () -> %                       empty? : % -> Boolean
--R eq? : (%,%) -> Boolean                hash : % -> SingleInteger
--R latex : % -> String                   listOfLists : % -> List(List(R))
--R map : ((R -> R),%) -> %               map : (((R,R) -> R),%,%) -> %
--R matrix : List(List(R)) -> %           maxColIndex : % -> Integer
--R maxRowIndex : % -> Integer            minColIndex : % -> Integer
--R minRowIndex : % -> Integer            ncols : % -> NonNegativeInteger
--R nrows : % -> NonNegativeInteger       one? : % -> Boolean
--R qelt : (%,Integer,Integer) -> R       recip : % -> Union(%,"failed")

--R retract : % -> R                      sample : () -> %
--R scalarMatrix : R -> %                 square? : % -> Boolean
--R squareMatrix : Matrix(R) -> %         symmetric? : % -> Boolean
--R trace : % -> R                        transpose : % -> %
--R zero? : % -> Boolean                  ?~=? : (%,%) -> Boolean
--R #? : % -> NonNegativeInteger if $ has finiteAggregate
--R ?*? : (DirectProduct(ndim,R),%) -> DirectProduct(ndim,R)
--R ?*? : (%,DirectProduct(ndim,R)) -> DirectProduct(ndim,R)
--R ?*? : (NonNegativeInteger,%) -> %
--R ?**? : (%,Integer) -> % if R has FIELD
--R ?**? : (%,NonNegativeInteger) -> %
--R ?/? : (%,R) -> % if R has FIELD
--R D : (%,NonNegativeInteger) -> % if R has DIFRING
--R D : (%,Symbol) -> % if R has PDRING(SYMBOL)
--R D : (%,List(Symbol)) -> % if R has PDRING(SYMBOL)
--R D : (%,Symbol,NonNegativeInteger) -> % if R has PDRING(SYMBOL)
--R D : (%,List(Symbol),List(NonNegativeInteger)) -> % if R has PDRING(SYMBOL)
--R D : (%,(R -> R),NonNegativeInteger) -> %
--R ?^? : (%,NonNegativeInteger) -> %
--R any? : ((R -> Boolean),%) -> Boolean if $ has finiteAggregate
--R characteristic : () -> NonNegativeInteger
--R coerce : Fraction(Integer) -> % if R has RETRACT(FRAC(INT))
--R column : (%,Integer) -> DirectProduct(ndim,R)
--R convert : % -> InputForm if R has KONVERT(INFORM)
--R count : (R,%) -> NonNegativeInteger if $ has finiteAggregate and R has SETCAT
--R count : ((R -> Boolean),%) -> NonNegativeInteger if $ has finiteAggregate
--R determinant : % -> R if R has commutative(*)
--R diagonal : % -> DirectProduct(ndim,R)
--R differentiate : % -> % if R has DIFRING
--R differentiate : (%,NonNegativeInteger) -> % if R has DIFRING
--R differentiate : (%,Symbol) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,List(Symbol)) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,Symbol,NonNegativeInteger) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,List(Symbol),List(NonNegativeInteger)) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,(R -> R),NonNegativeInteger) -> %
--R differentiate : (%,(R -> R)) -> %
--R eval : (%,List(R),List(R)) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,R,R) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,Equation(R)) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,List(Equation(R))) -> % if R has EVALAB(R) and R has SETCAT
--R every? : ((R -> Boolean),%) -> Boolean if $ has finiteAggregate
--R exquo : (%,R) -> Union(%,"failed") if R has INTDOM
--R inverse : % -> Union(%,"failed") if R has FIELD
--R less? : (%,NonNegativeInteger) -> Boolean
--R map! : ((R -> R),%) -> % if $ has shallowlyMutable
--R member? : (R,%) -> Boolean if $ has finiteAggregate and R has SETCAT
--R members : % -> List(R) if $ has finiteAggregate
--R minordet : % -> R if R has commutative(*)
--R more? : (%,NonNegativeInteger) -> Boolean
--R nullSpace : % -> List(DirectProduct(ndim,R)) if R has INTDOM

--R nullity : % -> NonNegativeInteger if R has INTDOM
--R parts : % -> List(R) if $ has finiteAggregate
--R rank : % -> NonNegativeInteger if R has INTDOM
--R reducedSystem : Matrix(%) -> Matrix(R)
--R reducedSystem : (Matrix(%),Vector(%)) -> Record(mat: Matrix(R),vec: Vector(R))
--R reducedSystem : (Matrix(%),Vector(%)) -> Record(mat: Matrix(Integer),vec: Vector(Integer)) if R has LINEXP(INT)
--R reducedSystem : Matrix(%) -> Matrix(Integer) if R has LINEXP(INT)
--R retract : % -> Fraction(Integer) if R has RETRACT(FRAC(INT))
--R retract : % -> Integer if R has RETRACT(INT)
--R retractIfCan : % -> Union(R,"failed")
--R retractIfCan : % -> Union(Fraction(Integer),"failed") if R has RETRACT(FRAC(INT))
--R retractIfCan : % -> Union(Integer,"failed") if R has RETRACT(INT)
--R row : (%,Integer) -> DirectProduct(ndim,R)
--R rowEchelon : % -> % if R has EUCDOM
--R size? : (%,NonNegativeInteger) -> Boolean
--R subtractIfCan : (%,%) -> Union(%,"failed")
--R
--E 36
--S 37 of 97
)set expose add constructor SQMATRIX
--I   SquareMatrix is now explicitly exposed in frame frame0
--E 37
--S 38 of 97
)show SQMATRIX
--R
--R SquareMatrix(ndim: NonNegativeInteger,R: Ring)  is a domain constructor
--R Abbreviation for SquareMatrix is SQMATRIX
--R This constructor is exposed in this frame.
--R Issue )edit bookvol10.3.pamphlet to see algebra source code for SQMATRIX
--R
--R------------------------------- Operations --------------------------------
--R ?*? : (R,%) -> %                      ?*? : (%,R) -> %
--R ?*? : (%,%) -> %                      ?*? : (Integer,%) -> %
--R ?*? : (PositiveInteger,%) -> %        ?**? : (%,PositiveInteger) -> %
--R ?+? : (%,%) -> %                      ?-? : (%,%) -> %
--R -? : % -> %                           ?=? : (%,%) -> Boolean
--R D : % -> % if R has DIFRING           D : (%,(R -> R)) -> %
--R 1 : () -> %                           0 : () -> %
--R ?^? : (%,PositiveInteger) -> %        antisymmetric? : % -> Boolean
--R coerce : % -> Matrix(R)               coerce : R -> %
--R coerce : Integer -> %                 coerce : % -> OutputForm
--R copy : % -> %                         diagonal? : % -> Boolean
--R diagonalMatrix : List(R) -> %         diagonalProduct : % -> R
--R elt : (%,Integer,Integer) -> R        elt : (%,Integer,Integer,R) -> R
--R empty : () -> %                       empty? : % -> Boolean
--R eq? : (%,%) -> Boolean                hash : % -> SingleInteger
--R latex : % -> String                   listOfLists : % -> List(List(R))
--R map : ((R -> R),%) -> %               map : (((R,R) -> R),%,%) -> %

--R matrix : List(List(R)) -> %           maxColIndex : % -> Integer
--R maxRowIndex : % -> Integer            minColIndex : % -> Integer
--R minRowIndex : % -> Integer            ncols : % -> NonNegativeInteger
--R nrows : % -> NonNegativeInteger       one? : % -> Boolean
--R qelt : (%,Integer,Integer) -> R       recip : % -> Union(%,"failed")
--R retract : % -> R                      sample : () -> %
--R scalarMatrix : R -> %                 square? : % -> Boolean
--R squareMatrix : Matrix(R) -> %         symmetric? : % -> Boolean
--R trace : % -> R                        transpose : % -> %
--R zero? : % -> Boolean                  ?~=? : (%,%) -> Boolean
--R #? : % -> NonNegativeInteger if $ has finiteAggregate
--R ?*? : (DirectProduct(ndim,R),%) -> DirectProduct(ndim,R)
--R ?*? : (%,DirectProduct(ndim,R)) -> DirectProduct(ndim,R)
--R ?*? : (NonNegativeInteger,%) -> %
--R ?**? : (%,Integer) -> % if R has FIELD
--R ?**? : (%,NonNegativeInteger) -> %
--R ?/? : (%,R) -> % if R has FIELD
--R D : (%,NonNegativeInteger) -> % if R has DIFRING
--R D : (%,Symbol) -> % if R has PDRING(SYMBOL)
--R D : (%,List(Symbol)) -> % if R has PDRING(SYMBOL)
--R D : (%,Symbol,NonNegativeInteger) -> % if R has PDRING(SYMBOL)
--R D : (%,List(Symbol),List(NonNegativeInteger)) -> % if R has PDRING(SYMBOL)
--R D : (%,(R -> R),NonNegativeInteger) -> %
--R ?^? : (%,NonNegativeInteger) -> %
--R any? : ((R -> Boolean),%) -> Boolean if $ has finiteAggregate
--R characteristic : () -> NonNegativeInteger
--R coerce : Fraction(Integer) -> % if R has RETRACT(FRAC(INT))
--R column : (%,Integer) -> DirectProduct(ndim,R)
--R convert : % -> InputForm if R has KONVERT(INFORM)
--R count : (R,%) -> NonNegativeInteger if $ has finiteAggregate and R has SETCAT
--R count : ((R -> Boolean),%) -> NonNegativeInteger if $ has finiteAggregate
--R determinant : % -> R if R has commutative(*)
--R diagonal : % -> DirectProduct(ndim,R)
--R differentiate : % -> % if R has DIFRING
--R differentiate : (%,NonNegativeInteger) -> % if R has DIFRING
--R differentiate : (%,Symbol) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,List(Symbol)) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,Symbol,NonNegativeInteger) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,List(Symbol),List(NonNegativeInteger)) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,(R -> R),NonNegativeInteger) -> %
--R differentiate : (%,(R -> R)) -> %
--R eval : (%,List(R),List(R)) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,R,R) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,Equation(R)) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,List(Equation(R))) -> % if R has EVALAB(R) and R has SETCAT
--R every? : ((R -> Boolean),%) -> Boolean if $ has finiteAggregate
--R exquo : (%,R) -> Union(%,"failed") if R has INTDOM
--R inverse : % -> Union(%,"failed") if R has FIELD
--R less? : (%,NonNegativeInteger) -> Boolean
--R map! : ((R -> R),%) -> % if $ has shallowlyMutable

--R member? : (R,%) -> Boolean if $ has finiteAggregate and R has SETCAT
--R members : % -> List(R) if $ has finiteAggregate
--R minordet : % -> R if R has commutative(*)
--R more? : (%,NonNegativeInteger) -> Boolean
--R nullSpace : % -> List(DirectProduct(ndim,R)) if R has INTDOM
--R nullity : % -> NonNegativeInteger if R has INTDOM
--R parts : % -> List(R) if $ has finiteAggregate
--R rank : % -> NonNegativeInteger if R has INTDOM
--R reducedSystem : Matrix(%) -> Matrix(R)
--R reducedSystem : (Matrix(%),Vector(%)) -> Record(mat: Matrix(R),vec: Vector(R))
--R reducedSystem : (Matrix(%),Vector(%)) -> Record(mat: Matrix(Integer),vec: Vector(Integer)) if R has LINEXP(INT)
--R reducedSystem : Matrix(%) -> Matrix(Integer) if R has LINEXP(INT)
--R retract : % -> Fraction(Integer) if R has RETRACT(FRAC(INT))
--R retract : % -> Integer if R has RETRACT(INT)
--R retractIfCan : % -> Union(R,"failed")
--R retractIfCan : % -> Union(Fraction(Integer),"failed") if R has RETRACT(FRAC(INT))
--R retractIfCan : % -> Union(Integer,"failed") if R has RETRACT(INT)
--R row : (%,Integer) -> DirectProduct(ndim,R)
--R rowEchelon : % -> % if R has EUCDOM
--R size? : (%,NonNegativeInteger) -> Boolean
--R subtractIfCan : (%,%) -> Union(%,"failed")
--R
--E 38
--S 39 of 97
)set expose add
--R----------------------------- The add Option ------------------------------
--R   The following groups are explicitly exposed in the current frame
--I      (called frame0 ):
--R                                   basic
--R                                categories
--R                                  naglink
--R                                   anna
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R                               SquareMatrix
--R
--R   When )set expose add is followed by no arguments, the information
--R      you now see is displayed. The arguments group and constructor are
--R      used to specify exposure groups or an explicit constructor to be
--R      added to the local frame exposure data. Issue
--R                            )set expose add group
--R      or
--R                         )set expose add constructor
--R      for more information.
--E 39
--S 40 of 97
)set expose drop constructor SQMATRIX

--I   SquareMatrix is now explicitly hidden in frame frame0
--E 40
--S 41 of 97
)show SQMATRIX
--R
--R SquareMatrix(ndim: NonNegativeInteger,R: Ring)  is a domain constructor
--R Abbreviation for SquareMatrix is SQMATRIX
--R This constructor is not exposed in this frame.
--R Issue )edit bookvol10.3.pamphlet to see algebra source code for SQMATRIX
--R
--R------------------------------- Operations --------------------------------
--R ?*? : (R,%) -> %                      ?*? : (%,R) -> %
--R ?*? : (%,%) -> %                      ?*? : (Integer,%) -> %
--R ?*? : (PositiveInteger,%) -> %        ?**? : (%,PositiveInteger) -> %
--R ?+? : (%,%) -> %                      ?-? : (%,%) -> %
--R -? : % -> %                           ?=? : (%,%) -> Boolean
--R D : % -> % if R has DIFRING           D : (%,(R -> R)) -> %
--R 1 : () -> %                           0 : () -> %
--R ?^? : (%,PositiveInteger) -> %        antisymmetric? : % -> Boolean
--R coerce : % -> Matrix(R)               coerce : R -> %
--R coerce : Integer -> %                 coerce : % -> OutputForm
--R copy : % -> %                         diagonal? : % -> Boolean
--R diagonalMatrix : List(R) -> %         diagonalProduct : % -> R
--R elt : (%,Integer,Integer) -> R        elt : (%,Integer,Integer,R) -> R
--R empty : () -> %                       empty? : % -> Boolean
--R eq? : (%,%) -> Boolean                hash : % -> SingleInteger
--R latex : % -> String                   listOfLists : % -> List(List(R))
--R map : ((R -> R),%) -> %               map : (((R,R) -> R),%,%) -> %
--R matrix : List(List(R)) -> %           maxColIndex : % -> Integer
--R maxRowIndex : % -> Integer            minColIndex : % -> Integer
--R minRowIndex : % -> Integer            ncols : % -> NonNegativeInteger
--R nrows : % -> NonNegativeInteger       one? : % -> Boolean
--R qelt : (%,Integer,Integer) -> R       recip : % -> Union(%,"failed")
--R retract : % -> R                      sample : () -> %
--R scalarMatrix : R -> %                 square? : % -> Boolean
--R squareMatrix : Matrix(R) -> %         symmetric? : % -> Boolean
--R trace : % -> R                        transpose : % -> %
--R zero? : % -> Boolean                  ?~=? : (%,%) -> Boolean
--R #? : % -> NonNegativeInteger if $ has finiteAggregate
--R ?*? : (DirectProduct(ndim,R),%) -> DirectProduct(ndim,R)
--R ?*? : (%,DirectProduct(ndim,R)) -> DirectProduct(ndim,R)
--R ?*? : (NonNegativeInteger,%) -> %
--R ?**? : (%,Integer) -> % if R has FIELD
--R ?**? : (%,NonNegativeInteger) -> %
--R ?/? : (%,R) -> % if R has FIELD
--R D : (%,NonNegativeInteger) -> % if R has DIFRING
--R D : (%,Symbol) -> % if R has PDRING(SYMBOL)
--R D : (%,List(Symbol)) -> % if R has PDRING(SYMBOL)
--R D : (%,Symbol,NonNegativeInteger) -> % if R has PDRING(SYMBOL)

--R D : (%,List(Symbol),List(NonNegativeInteger)) -> % if R has PDRING(SYMBOL)
--R D : (%,(R -> R),NonNegativeInteger) -> %
--R ?^? : (%,NonNegativeInteger) -> %
--R any? : ((R -> Boolean),%) -> Boolean if $ has finiteAggregate
--R characteristic : () -> NonNegativeInteger
--R coerce : Fraction(Integer) -> % if R has RETRACT(FRAC(INT))
--R column : (%,Integer) -> DirectProduct(ndim,R)
--R convert : % -> InputForm if R has KONVERT(INFORM)
--R count : (R,%) -> NonNegativeInteger if $ has finiteAggregate and R has SETCAT
--R count : ((R -> Boolean),%) -> NonNegativeInteger if $ has finiteAggregate
--R determinant : % -> R if R has commutative(*)
--R diagonal : % -> DirectProduct(ndim,R)
--R differentiate : % -> % if R has DIFRING
--R differentiate : (%,NonNegativeInteger) -> % if R has DIFRING
--R differentiate : (%,Symbol) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,List(Symbol)) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,Symbol,NonNegativeInteger) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,List(Symbol),List(NonNegativeInteger)) -> % if R has PDRING(SYMBOL)
--R differentiate : (%,(R -> R),NonNegativeInteger) -> %
--R differentiate : (%,(R -> R)) -> %
--R eval : (%,List(R),List(R)) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,R,R) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,Equation(R)) -> % if R has EVALAB(R) and R has SETCAT
--R eval : (%,List(Equation(R))) -> % if R has EVALAB(R) and R has SETCAT
--R every? : ((R -> Boolean),%) -> Boolean if $ has finiteAggregate
--R exquo : (%,R) -> Union(%,"failed") if R has INTDOM
--R inverse : % -> Union(%,"failed") if R has FIELD
--R less? : (%,NonNegativeInteger) -> Boolean
--R map! : ((R -> R),%) -> % if $ has shallowlyMutable
--R member? : (R,%) -> Boolean if $ has finiteAggregate and R has SETCAT
--R members : % -> List(R) if $ has finiteAggregate
--R minordet : % -> R if R has commutative(*)
--R more? : (%,NonNegativeInteger) -> Boolean
--R nullSpace : % -> List(DirectProduct(ndim,R)) if R has INTDOM
--R nullity : % -> NonNegativeInteger if R has INTDOM
--R parts : % -> List(R) if $ has finiteAggregate
--R rank : % -> NonNegativeInteger if R has INTDOM
--R reducedSystem : Matrix(%) -> Matrix(R)
--R reducedSystem : (Matrix(%),Vector(%)) -> Record(mat: Matrix(R),vec: Vector(R))
--R reducedSystem : (Matrix(%),Vector(%)) -> Record(mat: Matrix(Integer),vec: Vector(Integer)) if R has LINEXP(INT)
--R reducedSystem : Matrix(%) -> Matrix(Integer) if R has LINEXP(INT)
--R retract : % -> Fraction(Integer) if R has RETRACT(FRAC(INT))
--R retract : % -> Integer if R has RETRACT(INT)
--R retractIfCan : % -> Union(R,"failed")
--R retractIfCan : % -> Union(Fraction(Integer),"failed") if R has RETRACT(FRAC(INT))
--R retractIfCan : % -> Union(Integer,"failed") if R has RETRACT(INT)
--R row : (%,Integer) -> DirectProduct(ndim,R)
--R rowEchelon : % -> % if R has EUCDOM
--R size? : (%,NonNegativeInteger) -> Boolean
--R subtractIfCan : (%,%) -> Union(%,"failed")

--S 42 of 97
)set expose
--R
--R---------------------------- The expose Option ----------------------------
--R
--R Description: control interpreter constructor exposure
--R
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   basic
--R                                categories
--R                                  naglink
--R                                   anna
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   The following constructors are explicitly hidden in the current
--R      frame:
--R                               SquareMatrix
--R
--E 42
--S 43 of 97
)set expose drop group anna
--I   anna is no longer an exposure group for frame frame0
--E 43
--S 44 of 97
)set expose
--R
--R---------------------------- The expose Option ----------------------------
--R
--R Description: control interpreter constructor exposure
--R
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   basic
--R                                categories
--R                                  naglink
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   The following constructors are explicitly hidden in the current

--R      frame:
--R                               SquareMatrix
--R
--E 44
--S 45 of 97
)set expose add group
--R
--R---------------------------- The group Option -----------------------------
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   basic
--R                                categories
--R                                  naglink
--R
--Rbasic         naglink       anna          categories    Hidden
--Rdefaults
--E 45
--S 46 of 97
)set expose add group anna
--I   anna is now an exposure group for frame frame0
--E 46
--S 47 of 97
)set expose
--R
--R---------------------------- The expose Option ----------------------------
--R
--R Description: control interpreter constructor exposure
--R
--R   The following groups are explicitly exposed in the current frame
--R      (called initial ):
--R                                   anna
--R                                   basic
--R                                categories
--R                                  naglink
--R
--R   The following constructors are explicitly exposed in the current
--R      frame:
--R               there are no explicitly exposed constructors
--R
--R   The following constructors are explicitly hidden in the current
--R      frame:
--R                               SquareMatrix
--R
--E 47
--S 48 of 97
)display

--R
--R  )display keyword arguments are
--R     abbreviations
--R     all
--R     macros
--R     modes
--R     names
--R     operations
--R     properties
--R     types
--R     values
--R  or abbreviations thereof
--R
--E 48
--S 49 of 97
)display abb
--R
--R   You have requested that all abbreviations be displayed. As there are
--R      several hundred abbreviations, please confirm your request by
--R      typing y or yes and then pressing Enter :
--R
--R   >> System error:
--R   %.EOF is not of type SEQUENCE.
--R
--R   Continuing to read the file...
--R
--E 49
--S 50 of 97
)display all
--RProperties of %e :
--R   This is a system-defined macro.
--R   macro %e () == exp(1)
--RProperties of %i :
--R   This is a system-defined macro.
--R   macro %i () == complex(0,1)
--RProperties of %infinity :
--R   This is a system-defined macro.
--R   macro %infinity () == infinity()
--RProperties of %minusInfinity :
--R   This is a system-defined macro.
--R   macro %minusInfinity () == minusInfinity()
--RProperties of %pi :
--R   This is a system-defined macro.
--R   macro %pi () == pi()
--RProperties of %plusInfinity :
--R   This is a system-defined macro.
--R   macro %plusInfinity () == plusInfinity()
--RProperties of SF :

--R   This is a system-defined macro.
--R   macro SF () == DoubleFloat()
--E 50
--S 51 of 97
)display macros
--R
--RSystem-defined macros:
--R   macro %e () == exp(1)
--R   macro %i () == complex(0,1)
--R   macro %infinity () == infinity()
--R   macro %minusInfinity () == minusInfinity()
--R   macro %pi () == pi()
--R   macro %plusInfinity () == plusInfinity()
--R   macro SF () == DoubleFloat()
--E 51
--S 52 of 97
)display modes
--R   Type of value of %e:  (none)
--R   Type of value of %i:  (none)
--R   Type of value of %infinity:  (none)
--R   Type of value of %minusInfinity:  (none)
--R   Type of value of %pi:  (none)
--R   Type of value of %plusInfinity:  (none)
--R   Type of value of SF:  (none)
--E 52
--S 53 of 97
)display names
--R
--RNames of User-Defined Objects in the Workspace:
--R
--R   * None *
--R
--RNames of System-Defined Objects in the Workspace:
--R
--R%e                %i                %infinity         %minusInfinity
--R%pi               %plusInfinity     SF
--E 53
--S 54 of 97
)display operations
--R
--R   You have requested that all information about all AXIOM operations
--R      (functions) be displayed. As there are several hundred
--R      operations, please confirm your request by typing y or yes and
--R      then pressing Enter :
--R
--R   >> System error:

--R   %.EOF is not of type SEQUENCE.
--R
--R   Continuing to read the file...
--R
--E 54
--S 55 of 97
)display properties
--RProperties of %e :
--R   This is a system-defined macro.
--R   macro %e () == exp(1)
--RProperties of %i :
--R   This is a system-defined macro.
--R   macro %i () == complex(0,1)
--RProperties of %infinity :
--R   This is a system-defined macro.
--R   macro %infinity () == infinity()
--RProperties of %minusInfinity :
--R   This is a system-defined macro.
--R   macro %minusInfinity () == minusInfinity()
--RProperties of %pi :
--R   This is a system-defined macro.
--R   macro %pi () == pi()
--RProperties of %plusInfinity :
--R   This is a system-defined macro.
--R   macro %plusInfinity () == plusInfinity()
--RProperties of SF :
--R   This is a system-defined macro.
--R   macro SF () == DoubleFloat()
--E 55
--S 56 of 97
)display types
--R   Type of value of %e:  (none)
--R   Type of value of %i:  (none)
--R   Type of value of %infinity:  (none)
--R   Type of value of %minusInfinity:  (none)
--R   Type of value of %pi:  (none)
--R   Type of value of %plusInfinity:  (none)
--R   Type of value of SF:  (none)
--E 56
--S 57 of 97
)display values
--R   Value of %e:  (none)
--R   Value of %i:  (none)
--R   Value of %infinity:  (none)
--R   Value of %minusInfinity:  (none)
--R   Value of %pi:  (none)
--R   Value of %plusInfinity:  (none)

--R   Value of SF:  (none)
--E 57
--S 58 of 97
)display abb DHMATRIX
--R   DHMATRIX abbreviates domain DenavitHartenbergMatrix
--E 58
--S 59 of 97
)display abb DenavitHartenbergMatrix
--R   DHMATRIX abbreviates domain DenavitHartenbergMatrix
--E 59
--S 60 of 97
)display operations rotatex
--R
--R
--RThere is one exposed function called rotatex :
--R   [1] D1 -> DenavitHartenbergMatrix(D1) from DenavitHartenbergMatrix(
--R            D1)
--R            if D1 has Join(Field,TranscendentalFunctionCategory)
--R
--RExamples of rotatex from DenavitHartenbergMatrix
--R
--E 60
--S 61 of 97
)set fortran calling
--R                   Current Values of  calling  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rtempfile     set location of temporary data files       /tmp/
--Rdirectory    set location of generated FORTRAN files    ./
--Rlinker       linker arguments (e.g. libraries to search) -lxlf
--R
--E 61
--S 62 of 97
)set fortran calling tempfile
--R--------------------------- The tempfile Option ---------------------------
--R
--R Description: set location of temporary data files
--R
--R )set fortran calling tempfile  is used to tell AXIOM where
--R to place intermediate FORTRAN data files . This must be the
--R name of a valid existing directory to which you have permission
--R to write (including the final slash).
--R
--R Syntax:

--R   )set fortran calling tempfile DIRECTORYNAME
--R
--R The current setting is /tmp/
--E 62
--S 63 of 97
)set fortran calling tempfile /home/daly
--E 63
--S 64 of 97
)set fortran calling tempfile
--R--------------------------- The tempfile Option ---------------------------
--R
--R Description: set location of temporary data files
--R
--R )set fortran calling tempfile  is used to tell AXIOM where
--R to place intermediate FORTRAN data files . This must be the
--R name of a valid existing directory to which you have permission
--R to write (including the final slash).
--R
--R Syntax:
--R   )set fortran calling tempfile DIRECTORYNAME
--R
--R The current setting is /home/daly
--E 64
--S 65 of 97
)set fortran calling directory
--R-------------------------- The directory Option ---------------------------
--R
--R Description: set location of generated FORTRAN files
--R
--R )set fortran calling directory  is used to tell AXIOM where
--R to place generated FORTRAN files. This must be the name
--R of a valid existing directory to which you have permission
--R to write (including the final slash).
--R
--R Syntax:
--R   )set fortran calling directory DIRECTORYNAME
--R
--R The current setting is ./
--E 65
--S 66 of 97
)set fortran calling directory /home/daly/
--E 66
--S 67 of 97
)set fortran calling directory
--R-------------------------- The directory Option ---------------------------

--R
--R Description: set location of generated FORTRAN files
--R
--R )set fortran calling directory  is used to tell AXIOM where
--R to place generated FORTRAN files. This must be the name
--R of a valid existing directory to which you have permission
--R to write (including the final slash).
--R
--R Syntax:
--R   )set fortran calling directory DIRECTORYNAME
--R
--R The current setting is /home/daly/
--E 67
--S 68 of 97
)set fortran calling linker
--R---------------------------- The linker Option ----------------------------
--R
--R Description: linker arguments (e.g. libraries to search)
--R
--R )set fortran calling linkerargs  is used to pass arguments to the linker
--R when using  mkFort  to create functions which call Fortran code.
--R For example, it might give a list of libraries to be searched,
--R and their locations.
--R The string is passed verbatim, so must be the correct syntax for
--R the particular linker being used.
--R
--R Example: )set fortran calling linker "-lxlf"
--R
--R The current setting is -lxlf
--E 68
--S 69 of 97
)set fortran calling linker "-TPD"
--E 69
--S 70 of 97
)set fortran calling linker
--R---------------------------- The linker Option ----------------------------
--R
--R Description: linker arguments (e.g. libraries to search)
--R
--R )set fortran calling linkerargs  is used to pass arguments to the linker
--R when using  mkFort  to create functions which call Fortran code.
--R For example, it might give a list of libraries to be searched,
--R and their locations.
--R The string is passed verbatim, so must be the correct syntax for
--R the particular linker being used.
--R
--R Example: )set fortran calling linker "-lxlf"

--R
--R The current setting is -TPD
--E 70

———-

The kernel protect option has been disabled and will shortly be removed altogether. It was created because the CCL implementation hard coded some internal functions so they could not be changed.

— * —

--S 71 of 97 no longer supported
--)set kernel
--E 71
--S 72 of 97 no longer supported
--)set kernel warn
--E 72
--S 73 of 97 no longer supported
--)set kernel warn on
--E 73
--S 74 of 97 no longer supported
--)set kernel
--S 75 of 97 no longer supported
--)set kernel protect
--E 75
--S 76 of 97 no longer supported
--)set kernel protect on
--E 76
--S 77 of 97 no longer supported
--)set kernel
--E 77
--S 78 of 97
)set mes auto
--R
--R--------------------------- The autoload Option ---------------------------
--R
--R Description: print file auto-load messages
--R
--R The autoload option may be followed by any one of the following:
--R
--R    on
--R -> off

--R
--R The current setting is indicated.
--R
--E 78
--S 79 of 97
)set mes auto off
--E 79
--S 80 of 97
)set mes auto
--R
--R--------------------------- The autoload Option ---------------------------
--R
--R Description: print file auto-load messages
--R
--R The autoload option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 80
--S 81 of 97
)set mes auto on
--E 81
--S 82 of 97
)set mes auto
--R
--R--------------------------- The autoload Option ---------------------------
--R
--R Description: print file auto-load messages
--R
--R The autoload option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 82
--S 83 of 97
)lisp |$printLoadMsgs|
--R
--RValue = T
--E 83

--S 84 of 97
)set naglink
--R                   Current Values of  naglink  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------
--Rhost         internet address of host for NAGLink       localhost
--Rpersistence  number of (fortran) functions to remember  1
--Rmessages     show NAGLink messages                      on
--Rdouble       enforce DOUBLE PRECISION ASPs              on
--R
--E 84
--S 85 of 97
)set naglink host
--R----------------------------- The host Option -----------------------------
--R
--R Description: internet address of host for NAGLink
--R
--R )set naglink host is used to tell  AXIOM which  host to contact for
--R a NAGLink request. An Internet address should be supplied. The host
--R specified must be running the NAGLink daemon.
--R
--R The current setting is localhost
--E 85
--S 86 of 97
)set naglink persistence
--R------------------------- The persistence Option --------------------------
--R
--R Description: number of (fortran) functions to remember
--R
--R )set naglink persistence is used to tell  the  nagd  daemon how  many ASP
--R source and object files to keep around in case you reuse them. This helps
--R to avoid needless recompilations. The number specified should be a
--R non-negative integer.
--R
--R The current setting is 1
--E 86
--S 87 of 97
)set naglink messages
--R--------------------------- The messages Option ---------------------------
--R
--R Description: show NAGLink messages
--R
--R The messages option may be followed by any one of the following:
--R
--R -> on

--R    off
--R
--R The current setting is indicated.
--R
--E 87
--S 88 of 97
)set naglink double
--R---------------------------- The double Option ----------------------------
--R
--R Description: enforce DOUBLE PRECISION ASPs
--R
--R The double option may be followed by any one of the following:
--R
--R -> on
--R    off
--R
--R The current setting is indicated.
--R
--E 88
--S 89 of 97
)set naglink host axiom-developer.org
--E 89
--S 90 of 97
)set naglink host
--R----------------------------- The host Option -----------------------------
--R
--R Description: internet address of host for NAGLink
--R
--R )set naglink host is used to tell  AXIOM which  host to contact for
--R a NAGLink request. An Internet address should be supplied. The host
--R specified must be running the NAGLink daemon.
--R
--R The current setting is axiom-developer.org
--E 90
--S 91 of 97
)set naglink persistence 10
--E 91
--S 92 of 97
)set naglink persistence
--R------------------------- The persistence Option --------------------------
--R
--R Description: number of (fortran) functions to remember
--R
--R )set naglink persistence is used to tell  the  nagd  daemon how  many ASP
--R source and object files to keep around in case you reuse them. This helps

--R to avoid needless recompilations. The number specified should be a
--R non-negative integer.
--R
--R The current setting is 10
--E 92
--S 93 of 97
)set naglink messages off
--E 93
--S 94 of 97
)set naglink messages
--R--------------------------- The messages Option ---------------------------
--R
--R Description: show NAGLink messages
--R
--R The messages option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 94
--S 95 of 97
)set naglink double off
--E 95
--S 96 of 97
)set naglink double
--R---------------------------- The double Option ----------------------------
--R
--R Description: enforce DOUBLE PRECISION ASPs
--R
--R The double option may be followed by any one of the following:
--R
--R    on
--R -> off
--R
--R The current setting is indicated.
--R
--E 96
--S 97 of 97
)set naglink
--R                   Current Values of  naglink  Variables
--R
--RVariable     Description                                Current Value
--R-----------------------------------------------------------------------------

--Rhost         internet address of host for NAGLink       axiom-developer.org
--Rpersistence  number of (fortran) functions to remember  10
--Rmessages     show NAGLink messages                      off
--Rdouble       enforce DOUBLE PRECISION ASPs              off
--R
--E 97
)spool
)lisp (bye)

———-

References^{[1] nothing}

33